OpenStudy (anonymous):

when combining g(x) = 8x - 7 and f(x)= sq {x+3}, how do i combine them when x+3 is a square root? btw i have to find f(g(x)).

OpenStudy (anonymous):

what is \(f(x)\) ? is it \[f(x)=(x+3)^2\] ?

OpenStudy (anonymous):

\[\sqrt{x+3}\]

OpenStudy (anonymous):

oh square root

OpenStudy (anonymous):

\[f(g(x))=f(8x-7)=\sqrt{8x-7+3}\] then combine like terms, i.e. \(-7+3\)

OpenStudy (anonymous):

oooh alright